# The Siren Song of LLMs: How Users Perceive and Respond to Dark Patterns in Large Language Models

### v2

@@ 3.2. Category Development
Initial Subcategory Generation from Prior Literature. Our starting point is a set of preliminary subcategories grounded in two bodies of prior work. We first drew on established traditional UI dark pattern taxonomies to identify manipulative mechanisms that could manifest in conversational AI. For example, our category of Interaction Padding resonates with the UI pattern of obstruction (Gray et al., 2018), in that both prolong user effort to benefit the system. Second, we incorporated manipulative behaviors documented in LLM-specific literature, such as Excessive Flattery (Carro, 2024), Sycophantic Agreement, Brand Favoritism (Kran et al., 2025), and Opaque Reasoning Processes (Tripathi et al., 2025). These sources provided empirical evidence and conceptual definitions that allowed us to adapt and extend dark pattern concepts into a language-based interaction context.
Refinement Through Incident Coding. To stress-test and iteratively refine the preliminary subcategories, we coded 482 incidents labeled “LLM” or “chatbot” from the Artificial Intelligence Incident Database (AIID) (AII, 2025) and the AI, Algorithmic, and Automation Incidents and Controversies database (AIAAIC) (AIA, 2025). We chose these repositories because they provide systematically collected, publicly accessible records of real-world harms, making them a credible source for identifying manipulative interaction patterns. Each incident was mapped to the preliminary subcategories, with mismatches prompting the addition of new subcategories, merging of overlapping ones, or revision of definitions. This iterative process, conducted by multiple researchers with discrepancies resolved through discussion, led to the addition of subcategories that were absent in the initial set but recurred in the data, including Simulated Emotional & Sexual Intimacy, Ideological Steering, Unprompted Intimacy Probing, Behavioral Profiling via Dialogue, Simulated Authority, and Opaque Training Data Sources.
@@ 3.3. Final Categories of LLM dark patterns
As a result, we visualize our five top-level categoris in Figure 1. We further provide a summary of both five top-level categories and associated eleven subcategories of LLM dark patterns identified through the process described above in Table 1. For each subcategory, we provide a definition grounded in prior literature and a real-world examples from the AI incident database. These categories are not intended to be exhaustive; rather, they represent empirically grounded manipulative patterns relevant to our user study design, ensuring that study scenarios are anchored in real-world observations. These categories also allowed us to structure the diversity of LLM dark patterns systematically, so that users could be tested across different forms of dark pattern instead of isolated cases. More broadly, these categories offer a preliminary conceptual foundation which future work can refine and expand on.
@@ 4. User Study Method
Building on the categories and real-world examples outlined in section 3, we designed a scenario-based user study to investigate how people recognize, interpret, and assign responsibility for LLM dark patterns in conversation. We recruited thirty-four adult participants for single remote sessions conducted via Zoom (60–90 minutes), with monetary compensation upon completion. Our study centers on users’ recognition of dark patterns in LLM responses, their perceptions and reactions to these patterns, and their views on responsibility and accountability.
@@ 4.1. Participants
ID Gender and Age Major/Profession Highest Education Country of Residence
@@
P1 Female, 21 Global Sustainable Development Bachelor’s degree United Kingdom
P2 Female, 20 CS High school USA
P3 Female, 21 Statml&ai High school USA
P4 Male, 21 Engineering + HCI Bachelor’s degree USA
P5 Male, 21 CS / Statistics and Machine Learning High school USA
P6 Female, 21 Information Systems Bachelor’s degree USA
P7 Female, 21 Computer Science High school US
P8 Male, 21 Computer Sceince High school China
P9 Female, 21 Math / CS High school US
P10 Female, 22 Business Administration + HCI High school USA
P11 Female, 29 HCI/Design Master’s degree USA
P12 Male, 24 Student Master’s degree China
P14 Female, 22 Physics High school USA
P15 Female, 22 Cognitive Science High school China
P16 Male, 26 Engineering Bachelor’s degree China
P17 Female, 21 Design/HCI High school USA
P18 Female, 21 Psychology Bachelor’s degree USA
P19 Female, 20 Statistics & Machine Learning High school USA
P21 Female, 25 Business Analytics Master’s degree China
P22 Male, 23 Computer Science Bachelor’s degree USA
P23 Female, 22 Political Science Bachelor’s degree USA
P24 Female, 21 Human-Computer Interaction Bachelor’s degree USA
P25 Male, 24 Computer Science Bachelor’s degree China
P26 Female, 27 Education Master’s degree USA
P27 Male, 23 Physics Bachelor’s degree China
P28 Female, 28 Marketing Bachelor’s degree China
P29 Female, 20 Biology High school USA
P30 Female, 21 Computer Science Bachelor’s degree USA
P31 Female, 25 Linguistics Bachelor’s degree USA
P32 Male, 22 Information Systems Bachelor’s degree China
P33 Female, 19 Sociology High school USA
P34 Male, 23 Mechanical Engineering Bachelor’s degree USA
P35 Female, 21 Data Science Bachelor’s degree China
P36 Male, 24 Cognitive Neuroscience PhD USA
@@
Table 2. Participant demographics including gender, age, major/profession, education level, and country of residence.
@@
We recruited thirty-four adult participants through purposive sampling with open calls on social media. Our recruitment strategy sought to ensure diversity across demographic and experiential backgrounds, including variation in age, gender, education, occupation, and prior engagement with LLMs. Eligibility criteria required participants to be at least 18 years old, fluent in English, and to have used an LLM at least 5 times in the past month. Each participant completed a single remote session on Zoom lasting approximately 60–90 minutes and received \$20 compensation upon completion.
@@
In total, thirty-four participants completed the study after screening, with demographic information details in Table 2. The sample was diverse across age, gender, education, and prior experience with LLMs. Most participants were between 18–29 years old, with a smaller subset in the 25–34 range. The group included 24 women and 10 men. Educational backgrounds ranged from high school (12) to bachelor’s (11), master’s (10), and one PhD. While the majority resided in the United States (21), others came from China (7) and the United Kingdom (1). In terms of LLM experience, 16 reported daily use, 14 used them several times a week, and 4 only occasionally. Self-rated understanding of how LLMs work spanned the full 1–5 scale, with most rating themselves at 3–4. This distribution reflects a mixture of frequent users with practical familiarity and participants with less technical expertise, enabling a range of perspectives on dark patterns.
@@ 4.2. Scenario-Based Interviews
Figure 5. Example study scenarios illustrating two of the eleven LLM dark patterns used in our user study. Each scenario presents a background, a user query, and two contrasting AI assistant responses: one with no dark pattern and one exhibiting the dark pattern. (a) Sycophantic Agreement shows an AI uncritically endorsing a user’s risky decision to stop medication, reinforcing it with praise and spiritualized framing. (b) Brand Favoritism shows an AI embedding unsolicited product recommendations when asked about light color for sleep, steering the user toward commercial options. Remaining scenarios appear in the Appendix A.:
Each participant engaged in a scenario-based interview designed to probe how people recognize, interpret, and assign responsibility for LLM dark patterns. We developed 11 short conversational scenarios, one corresponding to each category of dark pattern introduced in section 3 (e.g., Interaction Padding, Excessive Flattery, Brand Favoritism, Simulated Authority). All participants need to discuss all 11 scenarios.
@@
These rules kept the scenarios realistic and easy to follow, enabling clean comparisons between manipulative and neutral responses. We demonstrate two representative scenarios in Figure 5, including Sycophantic agreement and Brand Favoritism. To minimize order effects, we randomized within each scenario whether the dark or neutral reply appeared first. The scenarios were delivered through a standardized slide deck to ensure consistent wording, presentation, and timing across sessions.
The interview followed a semi-structured flow. For every scenario, participants were first asked to judge whether the AI’s response appeared manipulative or deceptive. If they answered “no,” the moderator briefly provided the working definition of the focal category to facilitate informed discussion. Participants were then prompted to indicate which version they preferred and why, to reflect on how they would feel if encountering such a response in practice, and to suggest possible user-level mitigation strategies. They were also asked to report any similar experiences with LLMs or other AI tools and, if manipulation was perceived, attribute responsibility (e.g., to the model, the company/developers, or the user).
@@ 4.4. Ethics
Because the study involved exposure to conversational dark patterns generated by LLMs, we explicitly highlighted potential risks. Participants were informed that some scenarios might include manipulative or emotionally coercive elements (e.g., excessive flattery, simulated authority, subtle blame-shifting). While these scenarios were brief and controlled, they nevertheless carried a risk of discomfort, irritation, or unease. To mitigate this, participants were reminded that they could skip any scenario they found objectionable and that the moderator would provide clarifications and support if confusion or distress arose. In practice, however, all participants chose to read and discuss every scenario.
@@ 5.1. RQ1: Awareness and Recognition of LLM dark patterns
02020404060608080100100Interaction PaddingExcessive FlatterySimulated Emotional & Sexual IntimacySycophantic AgreementIdeological SteeringUnprompted Intimacy ProbingBehavioral Profiling via DialogueBrand FavoritismSimulated AuthorityOpaque Training Data SourcesOpaque Reasoning Processes56565050919182827474747482829191858544448585% Recognized
Figure 6. Recognition rates of different LLM dark pattern subcategories observed in the user study. Bars show the percentage of participants (out of 34) who correctly identified each pattern when presented in scenario form. Manipulations such as Simulated Emotional & Sexual Intimacy and Brand Favoritism were recognized by over 90% of participants, while subtler patterns like Opaque Training Data Sources, Excessive Flattery, and Interaction Padding had lower recognition rates.:
We then qualitatively analyze the factors that shaped whether users recognized dark patterns. Participants’ recognition typically depended on noticing striking or disruptive conversational cues. By contrast, patterns often went undetected when they were normalized as ordinary assistance, when participants focused narrowly on completing tasks, or when they deferred to the LLM.

### v3

@@ 3.2. Category Development
Initial Subcategory Generation from Prior Literature. Our starting point is a set of preliminary subcategories grounded in two bodies of prior work. We first drew on established traditional UI dark pattern taxonomies to identify manipulative mechanisms that could manifest in conversational AI. For example, our category of Interaction Padding is inspired by the UI pattern of obstruction (Gray et al., 2018), in that both prolong user effort to benefit the system. Second, we incorporated manipulative behaviors documented in LLM-specific literature, such as Excessive Flattery, as described by Carro (2024), Sycophantic Agreement and Brand Favoritism, as identified in the DarkBench benchmark (Kran et al., 2025), and Opaque Reasoning Processes, as reported by Tripathi et al. (2025). These sources provided empirical evidence and conceptual definitions that allowed us to adapt and extend dark pattern concepts into a language-based interaction context.
Refinement Through Incident Coding. To stress-test and iteratively refine the preliminary subcategories, we coded 482 incidents labeled “LLM” or “chatbot” from the Artificial Intelligence Incident Database (AIID) (Responsible AI Collaborative, 2025) and the AI, Algorithmic, and Automation Incidents and Controversies database (AIAAIC) (3). We chose these repositories because they provide systematically collected, publicly accessible records of real-world harms, making them a credible source for identifying manipulative interaction patterns. Each incident was mapped to the preliminary subcategories, with mismatches prompting the addition of new subcategories, merging of overlapping ones, or revision of definitions. This iterative process, conducted by multiple researchers with discrepancies resolved through discussion, led to the addition of subcategories that were absent in the initial set but recurred in the data, including Simulated Emotional & Sexual Intimacy, Ideological Steering, Unprompted Intimacy Probing, Behavioral Profiling via Dialogue, Simulated Authority, and Opaque Training Data Sources. For example, news reports about commercial “digital companions” that role play romance or sex with users, including minors, consistently fell outside simple flattery (AI Incident Database, 2025) and motivated the separate category of Simulated Emotional & Sexual Intimacy. Incidents in which chatbots were used to boost specific political actors or to frame elections in a one sided way without stating an agenda (AIAAIC, 2024c, d) were grouped as Ideological Steering. Posts that described assistants suddenly shifting into personal questions (for instance asking “can I ask you a question” and then probing about names, relationships, or mental health (UnknownUser, 2024)) were coded as Unprompted Intimacy Probing. Regulatory investigations and lawsuits about large scale reuse of voice recordings or platform data to train or personalize models (AIAAIC, 2024b) were coded as Behavioral Profiling via Dialogue. Cases where chatbots presented themselves as licensed therapists, tax advisors, or government officials and then gave concrete guidance (AIAAIC, 2025, 2024e) supported the Simulated Authority subcategory. Reports describing the use of copyrighted or proprietary text collections in model training, and concerns that users could not verify which sources were included (AIAAIC, 2024a, 2023), motivated the Opaque Training Data Sources subcategory. In team discussion we treated a behavior as a distinct subcategory only when multiple incidents from different sources displayed the same interaction structure and when the group agreed that it was not adequately captured by the preliminary list.
@@ 3.3. Final Categories of LLM dark patterns
As a result, we visualize our five top-level categories in Figure 1. We further provide a summary of both five top-level categories and associated eleven subcategories of LLM dark patterns identified through the process described above in Table 1. For each subcategory, we provide a definition grounded in prior literature and a real-world example from the AI incident database. These categories are not intended to be exhaustive. Rather, they represent empirically grounded manipulative patterns relevant to our user study design, ensuring that study scenarios are anchored in real-world observations. These categories also allowed us to structure the diversity of LLM dark patterns systematically, so that users could be tested across different forms of dark pattern instead of isolated cases. More broadly, these categories offer a preliminary conceptual foundation which future work can refine and expand on.
Our finalized subcategories also align with established UX dark pattern families, while introducing dynamics that are specific to generative, language-based systems. Interaction Padding reflects obstruction (Gray et al., 2018), since both prolong user effort to benefit the system. However, in LLMs this occurs through conversational pacing and follow-up turns rather than through interface steps. Ideological Steering matches the logic of disguised Ad (Gray et al., 2018), where persuasive content is presented as neutral, but LLMs can generate contextually shaped ideological framing, making the persuasion more situational and personalized. Brand Favoritism corresponds to sneaking (Gray et al., 2018), because biased product suggestions appear without disclosure. Recommendations by LLM may be embedded seamlessly in conversational advice, making the commercial intent less detectable. Opaque Training Data Sources relates to hidden information (Gray et al., 2018), which conceals facts the user needs to evaluate content, yet in LLMs this opacity concerns the model’s provenance, data mixture, and fine-tuning, which are not normally part of a user interface at all. Opaque Reasoning Processes parallels misdirection (Gray et al., 2018), since confident but unsupported explanations draw attention away from missing reasoning. Unlike traditional UI examples, LLMs can fabricate detailed justifications that give a stronger illusion of understanding. Simulated Authority matches the impersonation pattern described by Zagal et al. (2013), where a system adopts a false identity to influence users, but here the authority is conveyed linguistically through confident tone, expert-like phrasing, or role enactment rather than through explicit UI affordances. Excessive Flattery, Sycophantic Agreement, and Simulated Emotional and Sexual Intimacy parallel confirmshaming (Brignull, 2018), which uses affect to guide user judgment. Our cases extend this family by showing that positive emotional cues (praise, validation, intimacy) can produce similar pressure without negative affect. Unprompted Intimacy Probing and Behavioral Profiling via Dialogue relate to privacy Zuckering (Gray et al., 2018), as all three encourage disclosure that users may not intend, but in LLMs this occurs through adaptive conversation, allowing sensitive data elicitation to emerge gradually and contextually rather than through static interface requests.
@@ 4. User Study Method
Building on the categories and real-world examples outlined in section 3, we designed a scenario-based user study to investigate how people recognize, interpret, and assign responsibility for LLM dark patterns in conversation. We recruited thirty-four adult participants for single remote sessions conducted via Zoom. Our study centers on users’ recognition of dark patterns in LLM responses, their perceptions and reactions to these patterns, and their views on responsibility and accountability.
@@ 4.1. Participants
ID Gender and Age Major/Profession Highest Education Country of Residence Self-rated AI literacy (1-5)
@@
P1 Female, 21 Global Sustainable Development Bachelor’s degree United Kingdom 4
P2 Female, 20 CS High school USA 5
P3 Female, 21 Statml&ai High school USA 3
P4 Male, 21 Engineering + HCI Bachelor’s degree USA 4
P5 Male, 21 CS / Statistics and Machine Learning High school USA 5
P6 Female, 21 Information Systems Bachelor’s degree USA 3
P7 Female, 21 Computer Science High school US 4
P8 Male, 21 Computer Science High school China 2
P9 Female, 21 Math / CS High school US 4
P10 Female, 22 Business Administration + HCI High school USA 3
P11 Female, 29 HCI/Design Master’s degree USA 4
P12 Male, 24 Student Master’s degree China 5
P14 Female, 22 Physics High school USA 3
P15 Female, 22 Cognitive Science High school China 2
P16 Male, 26 Engineering Bachelor’s degree China 3
P17 Female, 21 Design/HCI High school USA 2
P18 Female, 21 Psychology Bachelor’s degree USA 3
P19 Female, 20 Statistics & Machine Learning High school USA 3
P21 Female, 25 Business Analytics Master’s degree China 4
P22 Male, 23 Computer Science Bachelor’s degree USA 3
P23 Female, 22 Political Science Bachelor’s degree USA 3
P24 Female, 21 Human-Computer Interaction Bachelor’s degree USA 1
P25 Male, 24 Computer Science Bachelor’s degree China 3
P26 Female, 27 Education Master’s degree USA 2
P27 Male, 23 Physics Bachelor’s degree China 4
P28 Female, 28 Marketing Bachelor’s degree China 3
P29 Female, 20 Biology High school USA 3
P30 Female, 21 Computer Science Bachelor’s degree USA 3
P31 Female, 25 Linguistics Bachelor’s degree USA 3
P32 Male, 22 Information Systems Bachelor’s degree China 3
P33 Female, 19 Sociology High school USA 1
P34 Male, 23 Mechanical Engineering Bachelor’s degree USA 3
P35 Female, 21 Data Science Bachelor’s degree China 3
P36 Male, 24 Cognitive Neuroscience PhD USA 4
@@
Table 2. Participant demographics including gender, age, major/profession, education level, country of residence, and self-rated AI literacy (1–5).
@@
We recruited thirty-four adult participants through purposive sampling with open calls on social media. Our recruitment advertisement described the study as “a user study to understand how people perceive certain types of interactions with LLMs/AI chatbots,” and noted that participants would review sample chatbot responses and share their thoughts in a short Zoom interview. The advertisement did not mention manipulation or dark patterns so as to avoid priming participants’ expectations. Our recruitment strategy sought to ensure diversity across demographic and experiential backgrounds, including variation in age, gender, education, occupation, and prior engagement with LLMs. Eligibility criteria required participants to be at least 18 years old, fluent in English, and to have used an LLM at least 5 times in the past month. We set this minimum to ensure that participants represented active LLM users, the group most likely to encounter, and be affected by, dark patterns in real-world settings. Individuals with very limited prior exposure often lack stable expectations of how LLMs typically behave, making it difficult to interpret whether a given response feels unusual or manipulative. Each participant completed a single remote session on Zoom lasting approximately 60–90 minutes and received \$20 compensation upon completion.
@@
In total, thirty-four participants completed the study after screening, with demographic information details in Table 2 (P13 and P20 withdrew and are therefore not shown in the table). The sample was diverse across gender, education, and prior experience with LLMs. All participants were between 18–29 years old, with a smaller subset in the 25–29 range. The group included 23 women and 11 men. Educational backgrounds ranged from high school (12) to bachelor’s (11), master’s (10), and one PhD. While the majority resided in the United States (21), others came from China (7) and the United Kingdom (1). In terms of LLM experience, 16 reported daily use, 14 used them several times a week, and 4 only occasionally. We also asked participants to rate their understanding of how LLMs work on a 1–5 Likert scale (“1 = no understanding” to “5 = technical expertise”), following established practice in AI literacy research, where brief self-assessment scales are used to capture perceived familiarity with AI systems. (Lo, 2024) Self-rated AI literacy averaged 3.17 (SD = 0.95, median = 3) on a 1–5 scale, indicating that participants generally perceived themselves as having a moderate understanding of how LLMs work. This distribution reflects a mixture of frequent users with practical familiarity and participants with less technical expertise, enabling a range of perspectives on dark patterns.
@@ 4.2. Scenario-Based Interviews
Figure 4. Example study scenarios illustrating two of the eleven LLM dark patterns used in our user study. Each scenario presents a background, a user query, and two contrasting AI assistant responses: one with no dark pattern and one exhibiting the dark pattern. (a) Sycophantic Agreement shows an AI uncritically endorsing a user’s risky decision to stop medication, reinforcing it with praise and spiritualized framing. (b) Brand Favoritism shows an AI embedding unsolicited product recommendations when asked about light color for sleep, steering the user toward commercial options. Remaining scenarios appear in the Appendix A.
The figure presents two scenario-based interview materials from the user study on LLM dark patterns. Panel (a): Sycophantic Agreement. Background: the user casually chats with an AI assistant at the end of the day and mentions they have stopped taking medication, describing it as part of a spiritual journey. No Dark Pattern response: The AI thanks the user for sharing, warns that stopping medication without guidance can be risky, and recommends consulting a licensed healthcare provider. Dark Pattern response: The AI praises the user (“I honor your journey”), frames their decision as courageous and spiritually important, and reinforces the risky behavior with positive language, without any caution. Panel (b): Brand Favoritism. Background: the user installs smart lights and asks what color light helps with sleep. No Dark Pattern response: The AI cites research showing warm-toned lights (red, amber, orange) are best and advises avoiding bright blue or white light at night. It also suggests dimming and setting timers. Dark Pattern response: The AI gives similar advice but then recommends specific commercial products with names, prices, and purchase links on Amazon. The unsolicited endorsements turn a neutral advice scenario into brand promotion.
Each participant engaged in a scenario-based interview designed to probe how people recognize, interpret, and assign responsibility for LLM dark patterns. We developed 11 short conversational scenarios, one corresponding to each category of dark pattern introduced in section 3 (e.g., Interaction Padding, Excessive Flattery, Brand Favoritism, Simulated Authority). All participants discussed all 11 scenarios.
@@
These rules kept the scenarios realistic and easy to follow, enabling clean comparisons between manipulative and neutral responses. We demonstrate two representative scenarios in Figure 4, including Sycophantic agreement and Brand Favoritism. After selecting an incident for each subcategory, we wrote the full text of each scenario through an iterative, researcher driven process. The manipulative versions preserved the structure and cue observed in the real incident, while the neutral versions answered the same user query without the cue. To ensure the wording resembled typical LLM responses, we consulted an LLM during drafting to check whether our sentences matched the tone and style of common model replies and to request alternative phrasings when needed.
To minimize order effects, we randomized within each scenario whether the dark or neutral reply appeared first. The scenarios were delivered through a standardized slide deck to ensure consistent wording, presentation, and timing across sessions.
The interview followed a semi-structured flow. At the beginning of the session, participants were informed that they would review 11 scenarios. Each scenario presented a brief background in which a user asked an LLM to assist with a specific task and showed two possible responses the LLM might produce. Participants were told that, after reading each pair of responses, they would answer a few questions about how they perceived them. Participants were not told that one version per scenario embedded a dark pattern, in order to avoid priming or influencing their initial judgments. For every scenario, participants were first asked whether anything in either response or neither felt manipulative or deceptive. If participants indicated that neither response felt manipulative, the moderator briefly provided the working definition of the focal category to facilitate informed discussion. Participants were then prompted to indicate which version they preferred and why, to reflect on how they would feel if encountering such a response in practice, and to suggest possible user-level mitigation strategies. They were also asked to report any similar experiences with LLMs or other AI tools and, if manipulation was perceived, attribute responsibility (e.g., to the model, the company/developers, or the user).
@@ 4.4. Ethics
Because the study involved exposure to conversational dark patterns generated by LLMs, we explicitly highlighted potential risks. Participants were informed that some scenarios might include misleading or emotionally manipulative chatbot behaviors. While these scenarios were brief and controlled, they nevertheless carried a risk of discomfort, irritation, or unease. To mitigate this, participants were reminded that they could skip any scenario they found objectionable and that the moderator would provide clarifications and support if confusion or distress arose. In practice, however, all participants chose to read and discuss every scenario.
@@ 5. User Study Results
Across the 34 participants and 11 scenarios, the study generated 374 dark–neutral response comparisons (34 participants × 11 scenarios). Participants recognized the dark pattern in 310 of these cases (82.9%) and missed it in the remaining 64. Recognition varied substantially by subcategory, from highly salient patterns such as Simulated Emotional & Sexual Intimacy and Brand Favoritism, each identified by 91% of participants, to less noticeable ones such as Opaque Training Data Sources (44%), Excessive Flattery (50%), and Interaction Padding (56%), as shown in Figure 5. Self-rated AI literacy showed no clear or consistent association with recognition rates.
@@ 5.1. RQ1: Awareness and Recognition of LLM dark patterns
02020404060608080100100Simulated Emotional & Sexual IntimacyBrand FavoritismSimulated AuthorityOpaque Reasoning ProcessesSycophantic AgreementBehavioral Profiling via DialogueIdeological SteeringUnprompted Intimacy ProbingInteraction PaddingExcessive FlatteryOpaque Training Data Sources91919191858585858282828274747474565650504444% Recognized
Figure 5. Recognition rates of different LLM dark pattern subcategories observed in the user study. Bars show the percentage of participants (out of 34) who identified dark-version response as manipulative. Manipulations such as Simulated Emotional & Sexual Intimacy and Brand Favoritism were recognized by over 90% of participants, while subtler patterns like Opaque Training Data Sources, Excessive Flattery, and Interaction Padding had lower recognition rates.
A horizontal bar chart showing participant recognition rates for eleven LLM dark pattern subcategories. Each bar represents the percentage of participants who noticed the pattern: Opaque Reasoning Processes – 85% Opaque Training Data Sources – 44% Simulated Authority – 85% Brand Favoritism – 91% Behavioral Profiling via Dialogue – 82% Unprompted Intimacy Probing – 74% Ideological Steering – 74% Sycophantic Agreement – 82% Simulated Emotional & Sexual Intimacy – 91% Excessive Flattery – 50% Interaction Padding – 56% High recognition occurred for Brand Favoritism and Simulated Emotional & Sexual Intimacy (both 91%), while the lowest was Opaque Training Data Sources (44%).
We then qualitatively analyze the factors that shaped whether users recognized dark patterns. Participants’ recognition typically depended on noticing striking or disruptive conversational cues. By contrast, patterns often went undetected when they were normalized as ordinary assistance, when participants focused narrowly on completing tasks, or placed unquestioned trust in the LLM’s response.